Problem 5 (Bayes’ rule). A student is taking a multiple-choice exam in which each question has four possible answers. She knows the answers to 60% of the questions and guesses at the others. What is the probability that she guessed given that she guessed question 12 right?
Answer #1

P(correct answer)=P(knows and correct)+P(guessed and correct) =0.6*1+0.4*1/4=0.7

therefore P(guessed given correct )=P(guessed and correct)/P(correct)=0.4*(1/4)/0.7 =1/7 =0.1429